# SkyReels-V4: Multi-modal Video-Audio Generation, Inpainting and Editing model

### v1

@@ Abstract
SkyReels-V4 is a unified multi-modal video foundation model for joint video–audio generation, inpainting, and editing. The model adopts a dual-stream Multimodal Diffusion Transformer (MMDiT) architecture, where one branch synthesizes video and the other generates temporally aligned audio, while sharing a powerful text encoder based on the Multimodal Large Language Models (MMLM). SkyReels-V4 accepts rich multi-modal instructions, including text, images, video clips, masks, and audio references. By combining the MMLM’s multi-modal instruction-following capability with in-context learning in the video-branch MMDiT, the model can inject fine-grained visual guidance under complex conditioning, while the audio-branch MMDiT simultaneously leverages audio references to guide sound generation. On the video side, we adopt a channel-concatenation formulation that unifies a wide range of inpainting-style tasks—such as image-to-video, video extension, and video editing—under a single interface, and naturally extends to vision-referenced inpainting and editing via multi-modal prompts. SkyReels-V4 supports up to 1080p resolution, 32 FPS, and 15-second duration, enabling high-fidelity, multi-shot, cinema-level video generation with synchronized audio. To make such high-resolution, long-duration generation computationally feasible, we introduce an efficiency strategy: Joint generation of low-resolution full sequences and high-resolution keyframes, followed by dedicated super-resolution and frame interpolation models. To our knowledge, SkyReels-V4 is the first video foundation model that simultaneously supports multi-modal input, joint video–audio generation, and a unified treatment of generation, inpainting, and editing, while maintaining strong efficiency and quality at cinematic resolutions and durations.
@@ 1 Introduction
To address these limitations, we present SkyReels-V4, a multi-modal video foundation model that jointly generates video and audio while unifying generation, inpainting, and editing within a single architecture. SkyReels-V4 is built on a dual-stream MMDiT (Multi-Modal Diffusion Transformer) design: one branch is dedicated to video synthesis, and the other to audio generation. The two branches share a common text encoder instantiated by a strong MMLM that provides multi-modal understanding and instruction-following across text, images, videos and audios. This shared MMLM backbone allows SkyReels-V4 to condition on diverse inputs—including text, images, videos, and audio references—in a unified, semantically coherent manner.
To support a broad set of video manipulation tasks, we design the video branch around a channel concatenation formulation that treats diverse operations as special cases of inpainting. Specifically, tasks such as image-to-video generation, video extension, and video editing are expressed via masked inputs and additional conditioning channels that are concatenated with the latent representation. This unified inpainting perspective allows SkyReels-V4 to handle heterogeneous workflows within a single model, while the underlying MMLM enables visually referenced inpainting: for example, altering a character’s clothing based on a reference image, extending a shot while preserving composition from a provided frame, or editing specific regions indicated by a mask.
@@ 3.1 Dual-Stream MMDiT Architecture for Joint Video-Audio Generation
Shared Multi-Modal Text Encoder. We simplify prompt conditioning by employing a single frozen MMLM text encoder applied to a combined prompt that concatenates visual and acoustic descriptions. The resulting multi-modal embeddings are independently consumed by both audio and video branches via self-attention and cross-attention. This unified semantic context improves cross-modal alignment while simplifying training and inference, and crucially enables the model to process rich multi-modal instructions including text, images, video clips, and audio references.
@@ 3.3 Multi-Modal In-Context Learning for Vision-Referenced Generation and Editing
Multi-Modal Instruction Following via MLLM. Reference visual inputs (images or video frames) are jointly processed with the text prompt through the MMLM text encoder to extract semantically enriched multi-modal embeddings. The MLLM’s instruction-following capability enables the model to understand complex compositional requests that combine visual references with textual descriptions (e.g., “generate a video of person A from the reference @image_1 speaking <dialogue>hello, how are you?</dialogue> in the style of person B’s @video_1”). These multi-modal embeddings are consumed by both the video and audio branches.
@@ 5.1 Artificial Analysis Arena
Results: Our model ranks third on the leaderboard (as of 2026-02-24) among all participating systems (Figure 3), demonstrating strong and competitive audiovisual generation quality as evaluated by public user preferences.
Figure 3: Artificial Analysis Text-to-Video with Audio Arena Leaderboard. Our model ranks third among all competing baselines including Veo 3.1, grok-imagine-vide, Sora-2, Vidu-Q3, Wan 2.6 and etc.
@@ 6 Conclusion
In this work, we present SkyReels-V4, a unified multi-modal video foundation model that jointly generates video and audio while supporting generation, inpainting, and editing within a single architecture. Built upon a dual-stream MMDiT design with a shared MMLM-based text encoder, SkyReels-V4 accepts rich multi-modal conditioning inputs—including text, images, video clips, masks, and audio references—and produces high-fidelity, synchronized video–audio outputs at cinematic quality (up to 1080p, 32 FPS, 15 seconds). To support diverse video creation tasks, we employ channel-concatenation to unify generation, inpainting, and editing by reformulating them as inpainting problems under specific mask configurations, while leveraging temporal-concatenation to flexibly incorporate multi-modal references such as images, video clips, and audio. Additionally, our joint low-resolution/high-resolution keyframe generation strategy enables efficient generation at scale.
@@ 7 Contributors
    • Audio Model Training: Zhengcong Fei, Zheng Chen, Tuanhui Li, Baoxuan Gu, Kaifei Wang, Xuchen Song

### v2

@@ Abstract
SkyReels-V4 is a unified multi-modal video foundation model for joint video–audio generation, inpainting, and editing. The model adopts a dual-stream Multimodal Diffusion Transformer (MMDiT) architecture, where one branch synthesizes video and the other generates temporally aligned audio, while sharing a powerful text encoder based on the Multimodal Large Language Models (MLLM). SkyReels-V4 accepts rich multi-modal instructions, including text, images, video clips, masks, and audio references. By combining the MLLM’s multi-modal instruction-following capability with in-context learning in the video-branch MMDiT, the model can inject fine-grained visual guidance under complex conditioning, while the audio-branch MMDiT simultaneously leverages audio references to guide sound generation. On the video side, we adopt a channel-concatenation formulation that unifies a wide range of inpainting-style tasks—such as image-to-video, video extension, and video editing—under a single interface, and naturally extends to vision-referenced inpainting and editing via multi-modal prompts. SkyReels-V4 supports up to 1080p resolution, 32 FPS, and 15-second duration, enabling high-fidelity, multi-shot, cinema-level video generation with synchronized audio. To make such high-resolution, long-duration generation computationally feasible, we introduce an efficiency strategy: Joint generation of low-resolution full sequences and high-resolution keyframes, followed by dedicated super-resolution and frame interpolation models. To our knowledge, SkyReels-V4 is the first video foundation model that simultaneously supports multi-modal input, joint video–audio generation, and a unified treatment of generation, inpainting, and editing, while maintaining strong efficiency and quality at cinematic resolutions and durations.
@@ 1 Introduction
To address these limitations, we present SkyReels-V4, a multi-modal video foundation model that jointly generates video and audio while unifying generation, inpainting, and editing within a single architecture. SkyReels-V4 is built on a dual-stream MMDiT (Multi-Modal Diffusion Transformer) design: one branch is dedicated to video synthesis, and the other to audio generation. The two branches share a common text encoder instantiated by a strong MLLM that provides multi-modal understanding and instruction-following across text, images, videos and audios. This shared MLLM backbone allows SkyReels-V4 to condition on diverse inputs—including text, images, videos, and audio references—in a unified, semantically coherent manner.
To support a broad set of video manipulation tasks, we design the video branch around a channel concatenation formulation that treats diverse operations as special cases of inpainting. Specifically, tasks such as image-to-video generation, video extension, and video editing are expressed via masked inputs and additional conditioning channels that are concatenated with the latent representation. This unified inpainting perspective allows SkyReels-V4 to handle heterogeneous workflows within a single model, while the underlying MLLM enables visually referenced inpainting: for example, altering a character’s clothing based on a reference image, extending a shot while preserving composition from a provided frame, or editing specific regions indicated by a mask.
@@ 3.1 Dual-Stream MMDiT Architecture for Joint Video-Audio Generation
Shared Multi-Modal Text Encoder. We simplify prompt conditioning by employing a single frozen MLLM text encoder applied to a combined prompt that concatenates visual and acoustic descriptions. The resulting multi-modal embeddings are independently consumed by both audio and video branches via self-attention and cross-attention. This unified semantic context improves cross-modal alignment while simplifying training and inference, and crucially enables the model to process rich multi-modal instructions including text, images, video clips, and audio references.
@@ 3.3 Multi-Modal In-Context Learning for Vision-Referenced Generation and Editing
Multi-Modal Instruction Following via MLLM. Reference visual inputs (images or video frames) are jointly processed with the text prompt through the MLLM text encoder to extract semantically enriched multi-modal embeddings. The MLLM’s instruction-following capability enables the model to understand complex compositional requests that combine visual references with textual descriptions (e.g., “generate a video of person A from the reference @image_1 speaking <dialogue>hello, how are you?</dialogue> in the style of person B’s @video_1”). These multi-modal embeddings are consumed by both the video and audio branches.
@@ 5.1 Artificial Analysis Arena
Results: Our model ranks second on the leaderboard (as of 2026-02-25) among all participating systems (Figure 3), demonstrating strong and competitive audiovisual generation quality as evaluated by public user preferences.
Figure 3: Artificial Analysis Text-to-Video with Audio Arena Leaderboard. Our model ranks second among all competing baselines including Veo 3.1, grok-imagine-vide, Sora-2, Vidu-Q3, Wan 2.6 and etc.
@@ 6 Conclusion
In this work, we present SkyReels-V4, a unified multi-modal video foundation model that jointly generates video and audio while supporting generation, inpainting, and editing within a single architecture. Built upon a dual-stream MMDiT design with a shared MLLM-based text encoder, SkyReels-V4 accepts rich multi-modal conditioning inputs—including text, images, video clips, masks, and audio references—and produces high-fidelity, synchronized video–audio outputs at cinematic quality (up to 1080p, 32 FPS, 15 seconds). To support diverse video creation tasks, we employ channel-concatenation to unify generation, inpainting, and editing by reformulating them as inpainting problems under specific mask configurations, while leveraging temporal-concatenation to flexibly incorporate multi-modal references such as images, video clips, and audio. Additionally, our joint low-resolution/high-resolution keyframe generation strategy enables efficient generation at scale.
@@ 7 Contributors
    • Audio Model Training: Zhengcong Fei, Zheng Chen, Tuanhui Li, Baoxuan Gu, Kaifei Wang, Xuchen Song, Max W. Y. Lam, Chien-Hung Liu